OpenStudy (henrietepurina):

Write x^2 - 2x - 3 = 0 in the form (x - a)2 = b, where a and b are integers.

OpenStudy (zzr0ck3r):

hint: \(x^2-2x-3 = (x-1)^2+-3-1\)

OpenStudy (zzr0ck3r):

this follow from \(x^2+bx+c = (x+\frac{b}{2})^2+c-(\frac{b}{2})^2\)